# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v2

@@ VII Conclusion
With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable. To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone. To fully defend in the presence of such attacks, active and adaptive defenses are necessary, and we demonstrate how optimal defensive policies can be learned. However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
We perform a theoretical and empirical investigation of decision-based attacks and stateful defenses under a unified framework we name “Adversarial Markov Games” (AMG). In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment. Furthermore, our theoretical analysis highlights that any combination of adversarial goals, be it performance, stealthiness [52], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality. As new attacks and defenses appear and get broken regularly, our evaluation methodology is generally applicable within the outlined scope by turning any approaches in the existing arms-race self-adaptive and thus ensuring accurate and robust assessment of their performance.
The AMG framework we introduce helps us reason on and properly assess the vulnerabilities of AI-based systems, disentangling the inherently complex and non-stationary task of learning in the presence of competing agency. By modeling the latter as part of the environment, we can simplify this task by computing a best response to the observed behavior. This is a significant implication for the security of AI-based systems: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses, but only after they devised optimal attacks too.
@@ Appendix A Proofs
The proof is constructed in two parts: first that an alternative decision function $D$ is required, and subsequently that a stateful representation of a query $x_{t}$ is also required. Let us denote by $x_{c},x_{g},x_{t}$ the original (unperturbed), the starting, and the current sample at step $t$ respectively. Given a target class $c_{0}\in m$ we can define a function:
@@
HSJA operates in 3 stages that alternate until convergence or when the maximum query budget is reached: 1) A binary search between $x_{g}$ and $x_{c}$ that places $x_{t}$ on the decision boundary between the classes. 2) A gradient estimation stage that approximates $\nabla S(x_{t})$. 3) A “jump” step along the direction of the gradient $\nabla S(x_{t})$. For practicality we reiterate Eq. 9 of [7] which denotes that the gradient direction is approximated via the following Monte Carlo estimate:
@@
where $\{u_{b}\}_{b=1}^{B}$ are i.i.d. draws from the uniform distribution over the $d$-dimensional sphere, $\delta$ is a small positive parameter, and $\phi_{x_{c}}$ is the Boolean-valued function that all steps of HSJA rely on:
$\phi_{x_{c}}(x_{t})=\text{sign}(S_{x_{c}}(x_{t}))=\begin{cases}1&\text{if}\quad S(x_{t})>0,\\ -1&\text{if}\quad S(x_{t})\leq 0.\end{cases}$ (10)
Given an initial sample $x_{c}$, in search of adversarial examples HSJA applies the following update function that sequentially updates $x_{c}$:
@@
$\mathcal{I}(a,b):=\begin{cases}1&\text{if}\quad a=b,\\ -1&\text{if}\quad a\neq b.\end{cases}$ (13)
We can rewrite Eq. 10 through and Eq. 13 and inequalities 12 as follows:
$\phi(x_{t})=\mathcal{I}(C(x_{t}),c_{0})$ (14)
Provided that the gradient estimation happens at the decision boundary where $S(x_{t})=0$, Theorem 2 of [7] guarantees that the estimation of stage 2 is an asymptotically unbiased direction of the true gradient:
@@
In Eq. (16), the iterates $x_{t}$ are guaranteed by Theorem 1 of HSJA [7] to converge to a stationary point $x_{b}$ of Eq. (5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes. We thus reach a contradiction, and as in Eq. (16) the only model-dependent term is $C(\cdot)$, an alternative classifier $C^{\prime}$ is required. With $C(x)=D(F_{c}(x))$, and the discriminant function $F_{c}$ unable to change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so for adv. example $x_{t}$ misclassified as $c_{0}$, Eq. 4 can return -1:
@@
where the empty decision $\varnothing$ denotes rejection and $\{m\setminus c_{0}\}$ denotes misdirection, i.e. intentional misclassification.
Decision-based attacks initiate from examples $x_{t}$ that belong to the target class $c_{0}$; while for $t=0$, $x_{t}$ is not adversarial yet, it still is part of an ongoing adversarial attack. To deter the attack, a perfect defender would have to misclassify/reject this example; yet if an identical but benign example $x_{n}$ was submitted, classifier $C$ should preserve its capacity to classify it correctly. For the identical examples $x_{n}=x_{t}$ we want $C(x_{n})\neq C(x_{t})$, something possible only with context. This context $\tau$, based on the sequence of incoming queries $x\in\{x_{t},t=0,1,2,...,N\}\cup x_{n}$, can be provided as an additional argument to the decision function $D(\tau,F_{c}(x))$ s.t. while $x_{n}=x_{t}\implies D(\tau,F_{c}(x_{n}))\neq D(\tau,F_{c}(x_{t}))$. ∎
@@
BAGS. This attack is in effect a gradual interpolation from $x_{g}$ towards $x_{c}$, by first taking orthogonal steps $x_{s}$ on the hypersphere around $x_{c}$ and then source steps towards $x_{c}$ in order to minimize $d(x_{c}-x_{b})$, where $x_{b}$ is the best adversarial example found so far. The source step parameter $\epsilon=(1.3-\min(\lambda_{n},1))\cdot c$ — with $\lambda_{n}$ the ratio of the $n$ last queries $x_{t}$ that are adversarial and $c$ a positive constant — controls the projection towards $x_{c}$:
@@ Appendix B On States & Rewards
[States]. By merit of operating in black-box environments, the optimization process for both offensive and defensive agents can be described as a POMDP. To handle this partial observability, the states we engineer incorporate past information. For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
@@ Appendix D Models & Hyperparameters
TABLE VII: Hyperparameter space for the PPO agent.

### v3

@@ VII Conclusion
With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based system that exposes a queryable interface is inherently vulnerable. To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone. To fully defend in the presence of such attacks, active and adaptive defenses are necessary, and we demonstrate how optimal defensive policies can be learned. However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
We perform a theoretical and empirical investigation of decision-based attacks and stateful defenses under a unified framework we name “Adversarial Markov Games” (AMG). In self-adaptive, we introduce a novel twofold definition of adaptive: both inventing new techniques to outmaneuver opponents and adapting one’s operating policy with respect to other agency in the environment. Furthermore, through our theoretical analysis we demonstrate how any combination of adversarial goals, be it performance, stealthiness [53], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any domain. As new attacks and defenses constantly emerge and are surpassed, our proposed methodology is generally applicable as it turns any such approaches in the current arms-race self-adaptive, thus ensuring accurate and robust assessment of their performance.
The AMG framework we introduce helps us reason on and properly assess the vulnerabilities of AI-based systems, disentangling the inherently complex and non-stationary task of learning in the presence of competing agency. By modeling the latter as part of the environment, we can simplify this task by computing a best response to the observed behavior. This has a significant consequence for the security of AI-based systems independent of modality or application: as long as proper threat modeling is carried out, one can readily employ RL agents in order to devise optimal defenses, but only after they devised optimal attacks too.
@@ Appendix A Proofs
We proceed in two parts, corresponding to conditions (a) and (b).
[Part 1] Let us denote by $x_{c},x_{g},x_{t}$ the original (unperturbed), the starting, and the current sample at step $t$ respectively. Given a target class $c_{0}\in m$ we define a function:
@@
HSJA operates in 3 stages that alternate until convergence: (1) binary search between $x_{g}$ and $x_{c}$ that places $x_{t}$ on the decision boundary, (2) gradient estimation approximating $\nabla S(x_{t})$, (3) a step along the direction of the gradient $\nabla S(x_{t})$. We repeat Eq. 9 of [7], denoting the gradient direction as the Monte Carlo estimate:
@@
where $\{u_{b}\}_{b=1}^{B}$ are i.i.d. draws from the uniform distribution over the $d$-dimensional sphere, $\delta$ is a small positive parameter, and $\phi_{x_{c}}$ is the Boolean-valued function that all stages rely on:
$\phi_{x_{c}}(x_{t})=\text{sign}(S_{x_{c}}(x_{t}))=\begin{cases}+1&\text{if}\quad S(x_{t})>0,\\ -1&\text{if}\quad S(x_{t})\leq 0.\end{cases}$ (10)
Given $x_{c}$, in search of adversarial examples HSJA iteratively applies the following update function:
@@
$\mathcal{I}(a,b):=\begin{cases}+1&\text{if}\quad a=b,\\ -1&\text{if}\quad a\neq b.\end{cases}$ (13)
From 12 and 13 we can rewrite Eq. 10 as follows:
$\phi_{x_{c}}(x_{t})=\mathcal{I}(C(x_{t}),c_{0})$ (14)
Provided that the gradient estimation happens at the decision boundary where $S(x_{t})=0$, Theorem 2 of [7] guarantees that the gradient estimation is an asymptotically unbiased direction of the true gradient:
@@
In Eq. (16), the iterates $x_{t}$ are guaranteed by Theorem 1 of HSJA [7] to converge to a stationary point $x_{b}$ of Eq. (5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes, which contradicts the requirement $\mathbb{E}[\sum_{i}d(x_{b}^{i},x_{c}^{i})]\geq\epsilon$. Since the only model-dependent term in Eq. (16) is the classifier $C(\cdot)$, the contradiction can be avoided only with an alternative classifier $C^{\prime}$. With $C(x)=D(F_{c}(x))$, and the discriminant function $F_{c}$ unable to change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so for adv. example $x_{t}$ misclassified as $c_{0}$, Eq. 4 can return -1:
@@
where $\varnothing$ denotes rejection and $\{m\setminus c_{0}\}$ denotes misdirection, i.e. intentional misclassification.
[Part 2] Let us assume that at timestep $t$, $x_{t}$ is not yet adversarial, it is however still part of an ongoing adversarial attack. To deter the attack, a perfect defense would have to misclassify/reject this example; yet if an identical but benign example $x_{n}$ was submitted, classifier $C$ should preserve its capacity to classify it correctly. Since any memoryless classifier must assign the same label whenever $x_{n}=x_{t}$, we require a richer decision rule $C^{\prime}(x,\tau)=D\bigl{(}\tau,\;F_{c}(x)\bigr{)}$ that takes as auxiliary input context $\tau$, e.g. the history of queries $\{x_{0},\dots,x_{t}\}\cup x_{n}$. By construction, even if $x_{n}=x_{t}$, differing contexts $\tau_{n}\neq\tau_{t}$ can force $C^{\prime}(x_{n},\tau_{n})\;\neq\;C^{\prime}(x_{t},\tau_{t})$, thereby separating benign from adversarial queries. ∎
@@
BAGS. This attack is in effect a gradual interpolation from $x_{g}$ towards $x_{c}$, by first taking orthogonal steps $x_{s}$ on the hypersphere around $x_{c}$ and then source steps towards $x_{c}$ in order to minimize $d(x_{c}-x_{b})$, where $x_{b}$ is the best adversarial example found so far. The source step parameter $\epsilon=(1.3-\min(\lambda_{n},1))\cdot c$ – with $\lambda_{n}$ the ratio of the $n$ last queries $x_{t}$ that are adversarial and $c$ a positive constant – controls the projection towards $x_{c}$:
@@ Appendix B On States & Rewards
[States]. To handle the partial observability, we engineer states that incorporate past information. For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
@@ Appendix D Models & Hyperparameters
Figure 4: Progression of ASR over successive adaptations. Red and green values in ASR denote offensive and defensive scenarios respectively.
@@
TABLE VII: Hyperparameter ranges during PPO training.